40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11929     Entry time: Tue Jan 12 19:38:31 2016     In reply to: 11926
Author: gautam 
Type: Update 
Category: LSC 
Subject: Frequently making noise 

EDITS 15Jan:

  1. Schematic of test setup added (Attachment #5). Note that the UGF measurements were made with the LPF and gain on the 'wrong' SR560, in a way defeating the purpose of having 2 SR560s in the setup. I only realised this after taking the measuements. But having done the loop algebra, I believe we can extract the necessary information, which is what has been done in subsequent plots...
  2. Koji pointed out that UGFs of ~100kHz was probably too high - this is when I took a closer look at the setup and realised the remarks made above in point 1. I realised we were in fact measuring the 'Process' open-loop TF. We can recover the loop TF by measuring the controller TF (which I did, see Attachment #3). The UGF for the PSL+X PLL loop is ~7.5kHz while that for PSL+Y is ~22kHz (both with a 1Hz LP on the SR560 and gain of x200).
  3. During the above investigations, I found that the measured TF for a 1Hz LP on the SR560 is weird - there seems to be a zero around 5kHz which gives some phase lead where one would expect a uniformly decaying gain and phase to be -90 degrees. Eric and I confirmed this behavioud on another SR560. Low-pass at 10kHz and high-pass at 1kHz seem to work fine. I will investigate this further when I get the time. Anyhow I don't think this affects anything as long as we measure the correct OLTF. It is still not clear to me why we even need this to lock the PLL...
  4. All the spectra (Attachment #4 and #5) are now calibrated taking into account the loop TF. I've added another panel with the spectra in V/rtHz as measured on the SR785, along with the SR560 output noise. I don't think any of the conclusions below are affected by these edits.

Summary:

I took several measurements today using the revised PLL scheme of using the Marconi just as an LO, and actuating on the Laser PZT to keep the PLL locked (I will put up a sketch soon). On the evidence of the attached plots (spectra of PLL control signal), I guess we can conclude the following:

  1. The AUX X laser's frequency noise performance is consistent with the levels expected from 'typical' NPRO numbers (and the datasheet), and is more or less consistent across different diode currents/crystal temperatures (? see below...).
  2. The diode current should be set to something less than 2.00 A
  3. Qualitatively, there is a difference in the shape of the spectra between the PSL+X and PSL+Y combinations above a couple of kHz. I don't know why we see this.

Attachment #2: Measured OLG of PLL for the PSL+X and PSL+Y combinations. The UGF in both cases looks to be above 100 kHz, so I didn't do any calibration for the spectra attached. The gain on the SR560 was set to 200 for all measurements.

Attachment #3: Measured spectra of PLL control signal for various diode currents, with one reading from the PSL+Y combination plotted for comparison. When we took some data last night, Eric noted that there was a factor of ~6 increase in the overall frequency spectrum level at higher currents, I will update the plots with last night's data as well shortly. I found it hardest to keep the PLL locked at a diode current of 2.00 A across all measurements.

Attachment #4: Measured spectra of PLL control signal at two different crystal temperatures. There does not seem to be any significant dependance on temperature, although I did only do the measurement at two temperatures.

Attachment #4 Attachment #1All the data used to make these plots (plus some that have yet to be added to the plots, I will update them).

Misc notes:

  • All measurements taken with two free-running lasers (PSL shutter closed)
  • The SR560 noise was measured with the input on the SR560 set to ground. 
  • In order to go from V/rtHz to Hz/rtHz on the plots, I used 1MHz/V for the X-end laser (which I verified by a quick measurement today to be approximately correct) and 4.6 MHz/V for the Y-end laser, based on an earlier measurement. 
  • I re-routed the long BNC cable to the Y-end, have yet to remove it. The BNC from the PDH setup at the X-end has been re-attached to the X-end NPRO.

Unrelated to this work:

When I came in this afternoon, I noticed that the PMC was unlocked. The usual procedure of turning the servo gain to -10dB and playing around with the DC output adjust slider on the MEDM screen did not work. Eric toggled a few buttons on the MEDM screen after which we were able to relock the PMC using the DC output adjust slider.

Attachment 1: 2016_01_AUXLaser.tar.gz  1007 kB
Attachment 2: OLGs.pdf  69 kB  Uploaded Fri Jan 15 22:42:31 2016  | Hide | Hide all | Show all
OLGs.pdf
Attachment 3: variedCurrent.pdf  146 kB  Uploaded Fri Jan 15 22:42:43 2016  | Show | Hide all | Show all
Attachment 4: variedTemp.pdf  132 kB  Uploaded Fri Jan 15 22:43:02 2016  | Hide | Hide all | Show all
variedTemp.pdf
Attachment 5: PLL_setup.pdf  310 kB  Uploaded Tue Jan 19 15:45:34 2016  | Hide | Hide all | Show all
PLL_setup.pdf
ELOG V3.1.3-